1.A 90.0-N box is pulled 10.0 m along a level surface by a rope. If the rope makes an angle of 20.0° with the surface, and the force exerted through the rope is 75.0 N, the work is done on the box is ___J. (Give your answer with the correct number of sig digs and do not include units)


2.A 25.0-N object is held 2.10 m above the ground. The potential energy of the object with respect to the ground is ___J. (Give your answer with the correct number of significant digits and do not include units).


1
Expert's answer
2021-07-19T09:49:21-0400

1)


"W=Fs\cos{20}\\W=(75)(10)\cos{20}=705\ J"

2)


"U=Wh\\U=(25)(2.1)=52.5\ J"